You throw an object in the air with 50 meters per second velocity and 30 degrees with respect to the horizontal direction. Calculate the displacement in the horizontal direction and how high the objects goes in the vertical direction. Assume that you climbed up a mountain that is 30 meters high and throw the object up.
Answer #1


C hina- t-->> 3om Uo= 50 mis ux = 50 cos 300 ml 2583 m/s My = 50 Sis 30° mis = 25 mis a = not unt + 12 axt? y = y + Myt + 1/2 t = 6:105 sec. - : Horizontal displacement is a= 0 + $513) (6.155) m. = 264-35 mi eg for y velocity is given by ty Vy = My t